OpenStudy (anonymous):

I am studying radical equations, how can I solve sqrt(x+16) - sqrt(x) = 2 I was shown that squaring each side of the equation will eliminate the radical, but even when I move one radical to the right side of the equation I end up cancelling x.

OpenStudy (anonymous):

\[\sqrt{x+16} - \sqrt(x) = 2\]

OpenStudy (anonymous):

Try squaring both sides twice.